If sinA=1/3, sinB=1/4 find sin(A+B), where A and B are acute

Grade:11

1 Answers

Arun
6 years ago
As we know
If sinA = 1/3
Then cos A = \sqrt8/3
Now
If sinB = ¼
Then cos B =\sqrt15/4
Hence
Sin(A+B) = sinA cosB + cos A sinB
Now put the values
Sin(A+B) = \sqrt15/12 + \sqrt8/12